# On the MST-ratio: Theoretical Bounds and Complexity of Finding the Maximum

### v2

@@ Abstract
Given a finite set of red and blue points in ${{\mathbb{R}}}^{d}$, the MST-ratio is defined as the total length of the Euclidean minimum spanning trees of the red points and the blue points, divided by the length of the Euclidean minimum spanning tree of their union. The maximum MST-ratio of a point set is the maximum MST-ratio over all non-trivial colorings of its points by red and blue. We prove that finding the maximum MST-ratio of a given point set is NP-hard when the dimension is part of the input. Moreover, we present a quadratic-time $3$-approximation algorithm for this problem. As part of the proof, we show that, in any metric space, the maximum MST-ratio is smaller than $3$. Additionally, we study the average MST-ratio over all colorings of a set of $n$ points. We show that this average is always at least $\frac{n-2}{n-1}$, and for $n$ random points uniformly distributed in a $d$-dimensional unit cube, the average tends to $d]{2}$ in expectation as $n$ approaches infinity.
@@
Minimum Spanning Tree, Discrete and Computational Geometry, NP-hardness, Approximation Algorithms
@@ 1 Introduction
Recently, motivated by applications in spatial biology, Cultrera et al. [8] studied the interactions between color classes in a colored point set from a topological point of view. They introduced the concept of MST-ratio as one of the measures for mingling of points with different colors in a colored point set. Investigating this measure opens the door to interesting discrete geometry questions related to Euclidean minimum spanning trees (EMST), which is the main focus of this paper. For a set $P$ of $n$ points in the Euclidean space ${{\mathbb{R}}}^{d}$, an Euclidean minimum spanning tree of $P$, denoted by $EMST(P)$, is a minimum spanning tree (MST) of the complete geometric graph on $P$ where the weight of each edge is the Euclidean distance between its endpoints. For a partition $P=R\cup B$ of $P$ into red and blue points, the EMST-ratio (sometimes referred to as the MST-ratio) of $P$ and $R$ is the total length of an Euclidean minimum spanning tree for each of $R$ and $B$, divided by the length of a Euclidean minimum spanning trees of $P$:
$\mu(P,R)=\frac{|EMST(R)|+|EMST(B)|}{|EMST(P)|}.$
@@
How much longer can EMSTs of two finite sets be compared to an EMST of their union? Given this union, $P$, we are interested in the maximum ratio, $\gamma(P)$, over all non-trivial partitioning of $P$ into two sets. Upper and lower bounds for $\gamma$ of certain classes of point sets can be found in [9, 13], see Section 1.1 for further details. However, limited research has been conducted on the maximum EMST-ratio in higher dimensional Euclidean spaces. Moreover, the question of whether there is an efficient algorithm to compute $\gamma$ for a given point set remains unanswered.
@@
The question can also be raised in the abstract setting. Namely, given a weighted graph $G$ with positive weights and a bi-partition $V(G)=R\cup B$ of the vertices of $G$, we define the MST-ratio of this partition as:
$\mu(G,R)=\frac{|MST(G[R])|+|MST(G[B])|}{|MST(G)|},$
@@
where $G[R]$ and $G[B]$ are the induced subgraphs of $G$ on the vertex sets $R$ and $B$, respectively, and $|H|$ is the total weight of the edges of the graph $H$. The maximum MST-ratio of $G$, denoted by $\gamma(G)$, is the maximum ratio over all non-trivial bi-partitions of $V(G)$.
@@
In the MAX-MST-ratio problem, we aim to find a partition $V(G)=R\cup B$ of the vertices of $G$ that maximizes the MST-ratio. As for the geometric setting, we introduce the MAX-EMST-ratio problem, in which given dimension $d$ and a finite point set $P$ in ${{\mathbb{R}}}^{d}$, we aim to find MAX-MST-ratio of the complete geometric graph on $P$, with Euclidean distances as weights of edges. When $d$ is a fixed number, we call the problem d-MAX-EMST-ratio (e.g. $2$-MAX-EMST-ratio, $3$-MAX-EMST-ratio).
@@
Our work establishes new bounds on the maximum and average MST-ratio. We also analyze the computational complexity of MAX-MST-ratio and study it through the lens of approximation algorithms. By an $\alpha$-approximation ($\alpha\geq 1$) for MAX-MST-ratio, we mean a polynomial time algorithm that for every weighted graph $G$ provides a bi-partition $V(G)=R\cup B$ satisfying $\alpha\cdot\mu(G,R)\geq\gamma(G)$.
@@
  • The MAX-MST-ratio problem is NP-hard and hard to approximate even within a factor of $O(n^{1-\varepsilon})$. (Theorems 2.1,2.7)
  • The MAX-EMST-ratio problem is NP-hard. (Theorem 3.4)
  • In any metric space the maximum MST-ratio is smaller than $3$. (Theorem 3.8)
  • There exists an ${O}(n^{2})$ time $3$-approximation algorithm for the MAX-EMST-ratio problem. (Theorem 3.10)
  • For every set of $n$ points in Euclidean space, the average EMST-ratio over all colorings is at least $\frac{n-2}{n-1}$. (Theorem 4.3)
  • The average EMST-ratio for a set of $n$ random points uniformly distributed in $[0,1]^{d}$ tends to $d]{2}$ as $n$ goes to infinity. (Theorem 4.1)
1.1 Related Works
The maximum EMST-ratio is closely related to the Steiner ratio of the Euclidean space. Recall that the Steiner tree of a point set is a tree that connects them via segments of minimum total length, allowing using some extra points. The Steiner ratio, $\rho_{d}$ of ${{\mathbb{R}}}^{d}$ is the infimum over all finite point sets in ${{\mathbb{R}}}^{d}$ of the length of the Steiner tree divided by the length of a minimum spanning tree of the set. Figure 1 shows an example of three points forming an equilateral triangle of side length $1$ with Steiner ratio $\sqrt{3}/2$. Gilbert and Pollack [15] conjectured that this is the best example in the plane and therefore $\rho_{2}=\sqrt{3}/2\approx 0.866$. The best-known lower bound for $\rho_{2}$ is $0.824$, due to Chung and Graham [7]. Important to us is also a universal lower bound on $\rho_{d}$ that works for every $d$. Gilbert and Pollack [15] presented a short proof for $\rho_{d}\geq 0.5$. Later, Graham and Hwang [16] showed $\rho_{d}\geq 0.577$ for any $d$.
Figure 1: The point set $P$ shown by square nodes is just the three vertices of an equilateral triangle of side length 1. The black edges form EMST(P) with length $2$. Adding an extra vertex in the center, the green edges form the minimum Steiner tree of $P$ with length $\sqrt{3}$.
@@ 1.1 Related Works
$|EMST(P)|/n^{1-1/d}\rightarrow{}\beta(d)$ (1)
with probability $1$, where $\beta(d)>0$ is a constant depending only on the dimension. Best known lower and upper bounds for $\beta(2)$ are approximately $0.6$ by Avram and Bertsimas [2], and $0.707$ by Gilbert [14], respectively.
@@ 1.2 Outline
Section 2 focuses on the abstract graphs and presents complexity results on the MAX-MST-ratio problem. Section 3 addresses the geometric setting, establishing bounds for the maximum MST-ratio and providing complexity results for MAX-EMST-ratio. Section 4 explores the average EMST-ratio across all colorings of a point set. Section 5 presents the experimental results regarding the maximum, average, and bipartite EMST-ratio for random point sets. Finally, Section 6 concludes with open questions to deepen understanding of the MST-ratio.
@@ 2 Maximum MST-ratio of Abstract Graphs
In this section, we study the complexity of finding the maximum MST-ratio for abstract graphs. As the primary result of this section, we show that MAX-MST-ratio is NP-hard. Throughout this section and the subsequent sections, we call an edges colorful, if the two endpoints of it have different colors, otherwise we call it monochromatic.
Theorem 2.1.
The problem MAX-MST-ratio is NP-hard.
Proof 2.2.
Let $\mathcal{I}$ be an instance of the well-known NP-complete $3$-SAT problem (see [10]), with a set of variables $X$ and set of clauses $\mathcal{C}$. Let $m=|\mathcal{C}|$ be the number of clauses of $\mathcal{I}$. W.l.o.g., we can assume that in $\mathcal{I}$ for each variable $x$ there exists a clause that contains $x$ and a clause that contains $\bar{x}$. Indeed, otherwise, one can deterministically decide the assignment of those variables and reach a smaller instance. Similarly, we can assume that in a fixed clause $C$, there is no variable $x$ that appears as $x$ and $\bar{x}$ as otherwise one can remove $C$ and again reach a smaller instance.
Using $\mathcal{I}$, we construct an instance $\mathcal{I^{\prime}}$ of MAX-MST-ratio on $14m+$ 1 vertices in the following way. For every clause $C$ we design the following clause gadget $H_{C}$:
  • We introduce a new vertex $C$.
  • For every variable $x\in X$ that belongs to $C$, we add a new vertex $x_{C}$.
  • For every variable $x\in X$ such that $\bar{x}$ is in $C$, we add a new vertex $\bar{x}_{C}$.
  • Let $S_{C}$ be the vertex set that contains $C$ and all the vertices $x_{C},\bar{x}_{C}$ for every variable $x\in X$. Note that $|S_{C}|=4$. For every pair of vertices $u,v\in S_{C}$, we set the weight of the edge $uv$ to $1$ (see Figure 2).
For every variable $x\in X$, we define the variable gadget for $x$ in which for every $C,C^{\prime}\in\mathcal{C}$, if $C$ contains $x$ and $C^{\prime}$ contains $\bar{x}$, we add an edge of weight $1$ from $x_{C}$ to $\bar{x}_{C^{\prime}}$ (see Figure 2). We also add a vertex $F$ in our construction. For every clause $C$ we set the weight of the edge $FC$ to $1$. All the other edges in this graph have a weight $2$.
$x_{C_{1}}$$\dots$$x_{C_{2}}$$\dots$$\dots$$x_{C_{r-1}}$$\dots$$x_{C_{r}}$$\dots$$\bar{x}_{{C^{\prime}}_{1}}$$\dots$$\bar{x}_{{C^{\prime}}_{2}}$$\dots$$\dots$$\bar{x}_{{C^{\prime}}_{t-1}}$$\dots$$\bar{x}_{{C^{\prime}}_{t}}$$\dots$(d)$C_{1}$$x_{C_{1}}$$\bar{y}_{C_{1}}$$z_{C_{1}}$(a)$C_{2}$$\bar{x}_{C_{2}}$$\bar{y}_{C_{2}}$$z_{C_{2}}$(b)$C_{3}$$\bar{x}_{C_{3}}$$\bar{y}_{C_{3}}$$\bar{z}_{C_{3}}$(c)
Figure 2: (a), (b), and (c) show the gadget for clauses $C_{1}=x\lor\bar{y}\lor z$, $C_{2}=\bar{x}\lor\bar{y}\lor z$, $C_{3}=\bar{x}\lor\bar{y}\lor\bar{z}$, respectively. (d) depicts the gadget of a variable $x\in X$. All the edges in the figure have a weight $1$.
For $\mathcal{I^{\prime}}$, we denote the maximum possible value of $|MST(R)|+|MST(B)|$ among all the possible colorings of vertices into sets $R$ and $B$ by $OPT(\mathcal{I^{\prime}})$. We remark that this is simply the product of the MST-ratio of $\mathcal{I^{\prime}}$ and the weight of an MST of $\mathcal{I^{\prime}}$. Let $H$ be the induced subgraph of $\mathcal{I^{\prime}}$ on the edges of weight $1$. Observe that $H$ contains $m$ vertex-disjoint cliques of size four which are simply the clause gadget $H_{C}$ for every $C\in\mathcal{C}$. We show that $\mathcal{I}$ is satisfiable if and only if $OPT(\mathcal{I^{\prime}})\geq 6m-2$.
Lemma 2.3.
If $\mathcal{I}$ is satisfiable, then $OPT(\mathcal{I^{\prime}})\geq 6m-2$.
Proof 2.4.
Let $A$ be an assignment that satisfies $\mathcal{I}$. We set the color of $F$ to blue and for every clause $C$ we set the color of the vertex $C$ to red. Now for every variable $x$:
  • if $A$ sets $x$ to true, then for every clause $C$ that contains $x$, we set the color of $x_{C}$ to blue and for every clause $C^{\prime}$ that contains $\bar{x}$, we set the color of $\bar{x}_{C^{\prime}}$ to red.
  • otherwise, for every clause $C$ that contains $x$, we set the color of $x_{C}$ to red and for every clause $C^{\prime}$ that contains $\bar{x}$, we set the color of $\bar{x}_{C^{\prime}}$ to blue.
Now we argue that if $R$ and $B$ are respectively the set of red vertices and the set of blue vertices, then $|MST(R)|+|MST(B)|\geq 6m-2$. Firstly, observe that by construction, if an edge of weight $1$ is monochromatic then its endpoints must belong to the same clause gadget. Therefore, the subgraph induced by red (and similarly blue) vertices can only have edges of weight $2$ and edges that belong to a clause gadget. Furthermore, since $A$ satisfies $\mathcal{I}$ then for every clause $C$ at least one of the vertices in $H_{C}$ is blue and by construction the vertex $C$ is colored in red. Consequently, at most two edges in $H_{C}$ belong to $MST(R)\cup MST(B)$. Therefore, $MST(R)\cup MST(B)$ contains at most $2m$ edges of weight $1$. Since $\mathcal{I}^{\prime}$ has $14m+$ 1 vertices, $MST(R)\cup MST(B)$ has $14m-$ 1 edges. Thus $|MST(R)|+|MST(B)|\geq 2m+2(4m-1-2m)=6m-2$, and hence the claim.
Lemma 2.5.
If $OPT(\mathcal{I^{\prime}})\geq 6m-2$, then $\mathcal{I}$ is satisfiable.
Proof 2.6.
Let $Q$ be a coloring of $V(\mathcal{I^{\prime}})$ by red and blue such that if $R$ and $B$ are the set of red and blue vertices then $MST(R)+MST(B)\geq 6m-2$. The monochromatic edges of weight $1$ do not contain a forest of size $12m+$ 1. Otherwise, since $MST(R)\cup MST(B)$ has precisely $14m-$ 1 edges, it would have at most $22m-$ 2 edges of weight $2$, so its weight would satisfy $|MST(R)|+|MST(B)|\leq(4m-1)+2m-2=6m-3$, which is a contradiction.
Assume that there is a monochromatic edge $uv$ of weight $1$ such that its endpoints do not belong to the same clause gadget. Let $E^{\prime}$ be a subset of edges that contains the edge $uv$ and exactly two monochromatic edges from each clause gadget. As clause gadgets are vertex-disjoint, the edges of $E^{\prime}$ induce a forest of size $12m+$ 1, a contradiction. This means that any monochromatic edge of weight $1$ belongs to a clause gadget.
Now we construct an assignment $A(Q)$, that satisfies $\mathcal{I}$. W.l.o.g., assume that $Q$ colors the vertex $F$ in blue. Then by the above argument for every $C\in\mathcal{C}$, the vertex $C$ has color red, and for every variable $x$, the color of the vertices $x_{C}$ and $\bar{x}_{C^{\prime}}$ are different in $Q$. Now consider a variable $x$. $A(Q)$ sets $x$ to true if and only if there exists a clause $C$ such that $Q$ colors $x_{C}$ by blue or it colors $\bar{x}_{C}$ by red. We claim that every clause in $\mathcal{I^{\prime}}$ is satisfied. Assume this is not the case and a clause $C$ is not satisfied. Then all the vertices in $H_{C}$ are colored in red. Let $E^{\prime}$ be a subset of edges that contains three edges from $H_{C}$ that yield a spanning tree and also for any other clause $C^{\prime}$, $E^{\prime}$ has two monochromatic edges. Again, $E^{\prime}$ induces a forest of size $12m+$ 1, a contradiction.
This completes the proof. We conclude that computing MAX-MST-ratio is NP-hard.
@@
In Section 3, we will use the NP-hardness of the MAX-MST-ratio to demonstrate that the geometric version of the problem (i.e., the MAX-EMST-ratio) is also NP-hard. Moreover, in the general setting, the MAX-MST-ratio problem is extremely difficult to approximate. Theorem 2.7 presents a strong inapproximability result for computing the MAX-MST-ratio.
Theorem 2.7.
@@
For every $0<\varepsilon\leq 1$, there is no polynomial time $O(n^{1-\varepsilon})$-approximation for MAX-MST-ratio problem, unless P=NP.
Proof 2.8.
@@
We prove that given an $\alpha$-approximation for MAX-MST-ratio, one can obtain a $2\alpha$-approximation for the MAX-Clique problem. Let $G=(V,E)$ be a graph on $n$ vertices on which we wish to find the maximum clique. From this, we create an instance of MAX-MST-ratio with the input graph $G^{\prime}$, where $G^{\prime}$ is a complete graph with $V(G^{\prime})=V$ and the weight of any edge $e\in E(G^{\prime})$ is $n$ if $e\in E$, and it is $1$ otherwise. Given a bi-partition of vertices into sets $R$ and $B$, we refer to $|MST(R)|+|MST(B)|$ as the value of this bi-partition.
@@
Now we show that given a coloring $Q$ of value $k$ for $G^{\prime}$, we can find a clique of size at least $\frac{\lfloor k/n\rfloor}{2}+1$ in $G$ in polynomial time. Let $R_{Q}$ and $B_{Q}$ denote a red MST and a blue MST of $G$, respectively, in the coloring $Q$. Given the weights in $G^{\prime}$ belong to $\{1,n\}$, and since $R_{Q}\cup B_{Q}$ has exactly $n-2$ edges, then $R_{Q}\cup B_{Q}$ has exactly $\lfloor k/n\rfloor$ edges of weight $n$. W.l.o.g, we can assume that $R_{Q}$ has at least $\frac{\lfloor k/n\rfloor}{2}$ edges of weight $n$. Now consider $F:=(V(R_{Q}),E^{\prime\prime})$, where $E^{\prime\prime}$ is the set of edges of weight $1$ in $R_{Q}$, and let $C_{1},\ldots,C_{r}$ be the connected components of $F$. Clearly, $r\geq\frac{\lfloor k/n\rfloor}{2}+1$. For every $1\leq i\leq r$, let $u_{i}$ be a vertex in $C_{i}$. As $R_{Q}$ is an MST on the set of red vertices, then for every $i<j$, the weight of the edge $u_{i}u_{j}$ in $G$ is $n$. Hence, $\{u_{1},u_{2},\ldots,u_{r}\}$ is a clique of size $r\geq\frac{\lfloor k/n\rfloor}{2}+1$ in $G$.

### v3

# 5600MB Eindhoven, the Netherlands

@@ Abstract
Given a finite set of red and blue points in ${{\mathbb{R}}}^{d}$, the MST-ratio is defined as the total length of the Euclidean minimum spanning trees of the red points and the blue points, divided by the length of the Euclidean minimum spanning tree of their union. The MST-ratio has recently gained attention due to its direct interpretation in topological models for studying point sets with applications in spatial biology. The maximum MST-ratio of a point set is the maximum MST-ratio over all proper colorings of its points by red and blue. We prove that finding the maximum MST-ratio of a given point set is NP-hard when the dimension is part of the input. Moreover, we present a quadratic-time $3$-approximation algorithm for this problem. As part of the proof, we show that, in any metric space, the maximum MST-ratio is smaller than $3$. Additionally, we study the average MST-ratio over all colorings of a set of $n$ points. We show that this average is always at least $\frac{n-2}{n-1}$, and for $n$ random points uniformly distributed in a $d$-dimensional unit cube, the average tends to $d]{2}$ in expectation as $n$ approaches infinity.
@@
Minimum Spanning Tree NP-hardness Discrete and Computational Geometry Approximation Algorithms
@@ 1 Introduction
Recently, motivated by applications in spatial biology, Cultrera et al. [9] studied the interactions between color classes in a colored point set from a topological point of view. To this end, they developed a framework based on the chromatic Delaunay mosaic and explored its combinatorial and topological properties; see also [4]. Moreover, they introduced the concept of MST-ratio as one of the measures for the mingling of points with different colors in a colored point set. Investigating this measure opens the door to interesting discrete geometry questions related to Euclidean minimum spanning trees (EMST), which is the main focus of this paper. For a set $P$ of $n$ points in the Euclidean space ${{\mathbb{R}}}^{d}$, an Euclidean minimum spanning tree of $P$, denoted by $\text{EMST}(P)$, is a minimum spanning tree (MST) of the complete geometric graph on $P$ where the weight of each edge is the Euclidean distance between its endpoints. For a partition $P=R\cup B$ of $P$ into red and blue points, the EMST-ratio (sometimes referred to as the MST-ratio) can be defined as follows.
$\mu(P,R)=\frac{|\text{EMST}(R)|+|\text{EMST}(P\setminus R)|}{|\text{EMST}(P)|}.$
@@
where $|\text{EMST}(X)|$ for a point set $X$ is the length of an Euclidean minimum spanning tree of $X$. How much longer can EMSTs of two finite sets be compared to an EMST of their union? Given a point set $P$, we are interested in the maximum ratio, $\gamma(P)$, over all proper partitionings of $P$ into two sets. The upper and lower bounds for $\gamma$ of certain classes of point sets can be found in [10, 14], see Section 1.1 for further details. However, limited research has been conducted on the maximum EMST-ratio in higher-dimensional Euclidean spaces. Moreover, the question of whether there is an efficient algorithm to compute $\gamma$ for a given point set remains unanswered.
@@
The question can also be raised in the abstract setting. Namely, given a weighted complete graph $G$ with positive weights and a bi-partition $V(G)=R\cup B$ of the vertices of $G$, we define the MST-ratio of this partition as:
$\mu(G,R)=\frac{|\text{MST}(G[R])|+|\text{MST}(G[P\setminus R])|}{|\text{MST}(G)|},$
@@
where $G[R]$ and $G[B]$ are the induced subgraphs of $G$ on the vertex sets $R$ and $B$, respectively, and $|H|$ is the total weight of the edges of the graph $H$. The maximum MST-ratio of $G$, denoted by $\gamma(G)$, is the maximum ratio over all proper bi-partitions of $V(G)$.
@@
In the MAX-MST-ratio problem, we aim to find a partition $V(G)=R\cup B$ of the vertices of $G$ that maximizes the MST-ratio. As for the geometric setting, we introduce the MAX-EMST-ratio problem, in which given dimension $d$ and a point set $P$ in ${{\mathbb{R}}}^{d}$, we aim to find the maximum EMST-ratio of $P$. When $d$ is a fixed number, we call the problem d-MAX-EMST-ratio.
@@
Our work establishes new bounds on the maximum and average MST-ratio. We also analyze the computational complexity of MAX-MST-ratio and study it through the lens of approximation algorithms. By an $\alpha$-approximation ($\alpha~{}\geq~{}1$) for MAX-MST-ratio, we mean a polynomial-time algorithm that for every weighted graph $G$ provides a bi-partition $V(G)=R\cup B$ satisfying $\alpha\cdot\mu(G,R)\geq\gamma(G)$.
@@
  • The MAX-MST-ratio problem is NP-hard and hard to approximate even within a factor of $O(n^{1-\varepsilon})$ (Theorems 2.1).
  • The MAX-EMST-ratio problem is NP-hard (Theorem 3.1).
  • In any metric space, the maximum MST-ratio is smaller than $3$ (Theorem 3.2).
  • There exists an ${O}(n^{2})$ time $3$-approximation algorithm for the MAX-EMST-ratio problem (Theorem 3.3).
  • For every set of $n$ points in Euclidean space, the average EMST-ratio over all colorings is at least $\frac{n-2}{n-1}$ (Theorem 4.2).
  • The average EMST-ratio for a set of $n$ random points uniformly distributed in $[0,1]^{d}$ tends to $d]{2}$ as $n$ goes to infinity (Theorem 4.1).
1.1 Related Work
The maximum EMST-ratio is closely related to the Steiner ratio of the Euclidean space. Recall that the Steiner tree of a point set is a tree that connects the points via segments of minimum total length, allowing using some extra points. The Steiner ratio, $\rho_{d}$ of ${{\mathbb{R}}}^{d}$, is the infimum over all finite point sets in ${{\mathbb{R}}}^{d}$ of the length of the Steiner tree divided by the length of an EMST of the set. Figure 1 shows an example of three points that form an equilateral triangle of side length $1$ with Steiner ratio $\sqrt{3}/2$. Gilbert and Pollack [16] conjectured that this is the most extreme example in the plane and therefore $\rho_{2}=\sqrt{3}/2\approx 0.866$. The best-known lower bound for $\rho_{2}$ is $0.824$, due to Chung and Graham [8]. It is also important for us to have a universal lower bound on $\rho_{d}$ that holds for all $d$. Gilbert and Pollack [16] presented a short proof for $\rho_{d}\geq 0.5$. Later, Graham and Hwang [17] showed $\rho_{d}\geq 0.577$ for any $d$.
Figure 1: The point set $P$ shown by square nodes is $3$ vertices of an equilateral triangle of side length $1$. The black edges form $\text{EMST}(P)$ with length $2$. Adding an extra vertex in the center, the green edges form the minimum Steiner tree of $P$ with length $\sqrt{3}$.
@@ 1.1 Related Work
$|\text{EMST}(P)|/n^{1-1/d}\rightarrow{}\beta(d)$ (1)
with probability $1$, where $\beta(d)>0$ is a constant depending only on the dimension. Best known lower and upper bounds for $\beta(2)$ are approximately $0.6$ by Avram and Bertsimas [2] and $0.707$ by Gilbert [15], respectively.
@@ 1.2 Outline
Section 2 focuses on abstract graphs and presents complexity results on the MAX-MST-ratio problem. Section 3 addresses the geometric setting, establishing bounds for the maximum MST-ratio and providing complexity results for MAX-EMST-ratio problem. Section 4 explores the average EMST-ratio across all colorings of a point set. Section 5 concludes with open questions to deepen understanding of the MST-ratio. Section 0.A.3 presents experimental results regarding the maximum, average, and bipartite EMST-ratio for random point sets.
@@ 2 Maximum MST-ratio of Abstract Graphs
In this section, we study the complexity of finding the maximum MST-ratio for abstract graphs. Throughout this section and the subsequent sections, we call an edge colorful if the two endpoints of it have different colors; otherwise, we call it monochromatic.
@@
As the main result of this section, we show that not only MAX-MST-ratio is NP-hard, but it is extremely hard to approximate. In Section 3, we will use the NP-hardness of the MAX-MST-ratio to demonstrate that the geometric version of the problem (i.e., the MAX-EMST-ratio) is also NP-hard. Theorem 2.1 presents a strong inapproximability result for the MAX-MST-ratio problem.
Theorem 2.1
@@
For every $0<\varepsilon\leq 1$, there is no polynomial-time $O(n^{1-\varepsilon})$-approximation with weights restricted to $1$ and $n$ for MAX-MST-ratio problem, unless P=NP.
Proof
@@
We prove that given an $\alpha$-approximation for MAX-MST-ratio, one can obtain a $2\alpha$-approximation for the MAX-Clique problem. Let $G=(V,E)$ be a graph on $n$ vertices on which we wish to find the maximum clique. From this, we create an instance of MAX-MST-ratio with the input graph $G^{\prime}$, where $G^{\prime}$ is a complete graph with $V(G^{\prime})=V$ and the weight of any edge $e\in E(G^{\prime})$ is $n$ if $e\in E$, and it is $1$ otherwise. Given a bi-partition of vertices into sets $R$ and $B$, we refer to $|\text{MST}(R)|+|\text{MST}(B)|$ as the value of this bi-partition.
@@
Conversely, given a coloring $Q$ of value $k$ for $G^{\prime}$, we show that we can find a clique of size at least $\frac{\lfloor k/n\rfloor}{2}+1$ in $G$ in polynomial time. Let $R_{Q}$ and $B_{Q}$ denote a red MST and a blue MST of $G$, respectively, in the coloring $Q$. Given that the weights in $G^{\prime}$ belong to $\{1,n\}$ and since $R_{Q}\cup B_{Q}$ has exactly $n-2$ edges, then $R_{Q}\cup B_{Q}$ has exactly $\lfloor k/n\rfloor$ edges of weight $n$. W.l.o.g, we can assume that $R_{Q}$ has at least $\frac{\lfloor k/n\rfloor}{2}$ edges of weight $n$. Consider $F:=(V(R_{Q}),E^{\prime\prime})$, where $E^{\prime\prime}$ is the set of edges of weight $1$ in $R_{Q}$. Let $C_{1},\ldots,C_{r}$ be the connected components of $F$. Clearly, we have $r\geq\frac{\lfloor k/n\rfloor}{2}+1$. For every $1\leq i\leq r$, let $u_{i}$ be a vertex in $C_{i}$. As $R_{Q}$ is an MST on the set of red vertices, then for every $i<j$, the weight of the edge $u_{i}u_{j}$ in $G$ is $n$. Hence, $\{u_{1},u_{2},\ldots,u_{r}\}$ is a clique of size $r\geq\frac{\lfloor k/n\rfloor}{2}+1$ in $G$.